THIS IS A CONVERSION FACTORS QUESTION

Camp nirvana sent a warning message over the CB radio. The glass skulls are patrolling the roadways and only allowing free passage with a payment of 100 kg glass. The latest report stated there will be a possible run in every 400 miles. Your team has some solid 30 lb glass blocks to use as bribe “money”. For a trip of 2695 miles, how many blocks should you take on your trip to ensure you have enough glass?

if you guys could answer my question that will be great please and thank you. :)
1
Expert's answer
2020-11-12T06:36:44-0500

First let's calculate how much bride will be required :

2965/400 = 6.7375


1kg = 2.2lb

100kg = 220 lb


Now calculate total amount of water : 220"\times" 6.7375 = 1482.25 lb


No. of blocks required = 1482.25/30 = 49.4 blocks

nearly equal to 50 glass blocks are required.